Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

coins

Expert replies
by shibal » Tue Jul 14, 2009 6:14 pm
i know it's easy, but i wanna see the different approaches (shortcuts) to the following question:

4 coins are tossed, what is the prob that at least 3 will come up heads??

oa [spoiler]5/16[/spoiler]
Join the discussion
Source: — Problem Solving |

by pops » Tue Jul 14, 2009 9:54 pm
Probability that all 4 shows heads: 1/2*1/2*1/2*1/2 = 1/16
Probability that 3 shows heads: 4C1*1/2*1/2*1/2*1/2 = 4/16
probability at least 3 shows heads=1/16+4/16=5/16
Join the discussion

by tohellandback » Tue Jul 14, 2009 10:22 pm
i would like to solve it by taking that opposite. i.e. by taking the probability of not getting at least 3 heads.

Can anyone help??
The powers of two are bloody impolite!!
Join the discussion

by nitya34 » Tue Jul 14, 2009 10:37 pm
is the OA ok?
I believe if we use Bionomial Theorem

we get

(4C3) [(1/2)^3][(1/2)^1]=4(1/2)^4=4/16

https://gwydir.demon.co.uk/jo/probability/info.htm
Many of the great achievements of the world were accomplished by tired and discouraged men who kept on working.
Join the discussion

by scoobydooby » Tue Jul 14, 2009 10:51 pm
prob of atleast 3 coins showing head:

(prob of 3 coins head, 1 coin tail)+(prob of all coins head)

nitya34, you are missing out on another possibilty- all coins showing head.
so if you add 4C4*(1/2)^4=(1/2)^4=1/16 to the 4/16 you already have, you end up with 5/16
Join the discussion

by scoobydooby » Tue Jul 14, 2009 10:58 pm
tohellandback wrote:i would like to solve it by taking that opposite. i.e. by taking the probability of not getting at least 3 heads.

Can anyone help??
prob of getting atleast 3 heads= 1- prob of not getting atleast 3 heads.

prob of not getting atleast 3 heads= (prob of no head)+ (prob of 1 head)+ (prob of 2 head)

=4C4*(1/2)^4+4C1*(1/2)^4+4C2*(1/2)^4
=1/16+4/16+6/16
=11/16

so prob of getting atleast 3 heads=1-11/16=5/16
Join the discussion

Re: coins

by Stuart@KaplanGMAT » Tue Jul 14, 2009 11:07 pm
shibal wrote:i know it's easy, but i wanna see the different approaches (shortcuts) to the following question:

4 coins are tossed, what is the prob that at least 3 will come up heads??

oa [spoiler]5/16[/spoiler]
Super fast way to answer coin flip questions - use Pascal's Triangle.

The n=4 row of the triangle is 1 4 6 4 1.

We want at least 3 heads, which means we add up the last two numbers of the row.. 4+1=5.. then we divide that by the sum of the row... 16.

Answer: 5/16

For more on the triangle (and other coin flip tips), take a look at https://www.beatthegmat.com/coin-flip-qu ... 17911.html
Image

Stuart Kovinsky | Kaplan GMAT Faculty | Toronto

Kaplan Exclusive: The Official Test Day Experience | Ready to Take a Free Practice Test? | Kaplan/Beat the GMAT Member Discount
BTG100 for $100 off a full course
Join the discussion